OpenStudy (ray1998):

Please help! Will give both fan and medal :) Divide and simplify. 7 sqrt 14x^6 / 7 sqrt 2x^5 Thank you in advance!

jimthompson5910 (jim_thompson5910):

Is it this? \[\Large \frac{7\sqrt{14x^6}}{7\sqrt{2x^5}}\] or is it this \[\Large \frac{\sqrt[7]{14x^6}}{\sqrt[7]{2x^5}}\]

OpenStudy (ray1998):

The first :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok so the first thing we do is we can cancel out the 7s since 7 over 7 = 1 \[\Large \frac{7\sqrt{14x^6}}{7\sqrt{2x^5}}\] \[\Large \frac{\color{red}{7}\sqrt{14x^6}}{\color{red}{7}\sqrt{2x^5}}\] \[\Large \frac{\color{red}{\cancel{7}}\sqrt{14x^6}}{\color{red}{\cancel{7}}\sqrt{2x^5}}\] \[\Large \frac{\sqrt{14x^6}}{\sqrt{2x^5}}\] agreed?

OpenStudy (ray1998):

Agreed

jimthompson5910 (jim_thompson5910):

then we use the rule \[\Large \frac{\sqrt{x}}{\sqrt{y}} = \sqrt{\frac{x}{y}}\] to get \[\Large \frac{\sqrt{14x^6}}{\sqrt{2x^5}}=\sqrt{\frac{14x^6}{2x^5}}\] what comes next?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

this is a step by step sort of a thing

OpenStudy (ray1998):

Divide?

OpenStudy (ray1998):

You're fine! @Houdini_Dragon yes, it is :)

jimthompson5910 (jim_thompson5910):

yes, you need to simplify \(\LARGE \frac{14x^6}{2x^5}\)

OpenStudy (ray1998):

Sorry! My internet messed up. Is it, 7x^11?

jimthompson5910 (jim_thompson5910):

when you divide variables like that, you SUBTRACT the exponents

jimthompson5910 (jim_thompson5910):

Rule: \[\LARGE \frac{x^a}{x^b} = x^{a-b}\]

jimthompson5910 (jim_thompson5910):

Example \[\LARGE \frac{x^9}{x^4} = x^{9-4}=x^5\]

jimthompson5910 (jim_thompson5910):

`Is it, 7x^11?` close but not quite 100% correct

OpenStudy (ray1998):

7x^-1 ?

OpenStudy (ray1998):

@jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

the exponent up top is 6 the exponent in the bottom is 5 (top exponent) - (bottom exponent) = 6-5 = 1 so 1 is the final exponent

jimthompson5910 (jim_thompson5910):

\[\LARGE \frac{x^6}{x^5} = x^{6-5}=x^1 = x\]

OpenStudy (ray1998):

Yeh, I noticed I made that mistake, I meant positive 1, not negative.

jimthompson5910 (jim_thompson5910):

So all this means \[\Large \sqrt{\frac{14x^6}{2x^5}} = \sqrt{7x}\]

OpenStudy (ray1998):

Ahh, I see. And that's our final answer, correct?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (ray1998):

Okay. Thank you very much :)

jimthompson5910 (jim_thompson5910):

\[\Large \frac{7\sqrt{14x^6}}{7\sqrt{2x^5}}=\sqrt{7x}\]

jimthompson5910 (jim_thompson5910):

np